# Aligning Transformers with Continuous Feedback via Energy Rank Alignment

### v2

@@ 2 Energy rank alignment
3 Experiments
3.1 Generating small molecules with desired properties
We use a decoder-only representation for the molecular generator [4], where the generator has 2 layers, an embedding dimension of 512, a vocabulary of 324 tokens, and totals 3.5M parameters. Starting from a random initialization, we carry out pretraining on a dataset of 2.4M small molecules from the ChEMBL database [46] for 180 epochs. This version of the model is not conditioned on a prompt and generates a small molecule (represented as a SMILES string) given just a start-of-sequence token.
Central to ERA is, of course, access to a computable energy function. As a proof-of-concept, we first consider 5 different properties for which the corresponding energy function is easily evaluable: Quantitative Estimate of Drug-Likeness (QED) [6], Wildman-Crippen LogP (LogP) [43], Ring Count, Molar Refractivity (MR) [43], and Tanimoto Similarity [30] (Section 3.1.13.1.2). Briefly, LogP is a measure of the hydrophobicity of a molecule, MR is a measure of the polarizability of the molecule, and Tanimoto similarity is a measure of the similarity between two molecules (see Appendix E.2). We then investigate ERA in a more challenging context: generating small-molecules with high predicted bioactivity for two kinases (Section 3.1.3).
3.1.1 Unprompted molecular alignment on RDKit oracles
First, we independently target four different properties using ERA with an unprompted molecular generator (Fig. 2). Using the pretrained model as our reference policy, we generate a dataset $\mathcal{D}=\{\boldsymbol{y}_{1}^{(i)},\boldsymbol{y}_{2}^{(i)},U(\boldsymbol{y}_{1}^{(i)}),U(\boldsymbol{y}_{2}^{(i)})\}_{i=1}^{N}$ and carry out energy rank alignment on $\pi_{\boldsymbol{\theta}},$ where $\pi_{\boldsymbol{\theta}}$ is initialized using the weights of $\pi_{\rm ref}.$ Here, $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\sim\pi_{\rm ref}$ and $\boldsymbol{y}$ and $U(\boldsymbol{y})$ denote the generated molecule and its corresponding energy, respectively. For MR, Ring Count, and LogP, we define the energy $U$ to be a harmonic potential centered at a target value. For QED, we define the energy to be the negative logarithm of QED and vary $\beta$ to assess its impact on alignment (see Tables S2S3). In Fig. 2, we see that we successfully shift the distribution to target means that are both greater and lower than the average value of MR, Ring Count, and LogP under the reference policy. Furthermore, in the alignment of QED, we observe the effect of changing $\beta$ on the learned policy; with increased $\beta,$ the learned policy concentrates around low-energy samples (i.e. near $\textrm{QED}=1$), and with lower $\beta,$ the learned policy samples a greater range of QED values, as expected. We note that for each of these four experiments, we did not regularize towards the reference policy (i.e. $\gamma=0$). Even so, we were able to maintain both sample diversity and maintain appreciable sample validity (see Fig. S8 and Table S4).
Many molecular design tasks require balancing multiple properties, and designing an objective for multi-property alignment is straightforward within the ERA framework. To demonstrate this, we generate molecules with both high QED and LogP using ERA with an energy function weighted by property-specific $\beta$: $U=\beta_{\rm QED}U_{\rm QED}+\beta_{\rm LogP}U_{\rm LogP}$ (see Tables S2S5 for details on the energy function). We carry out ERA with different pairs of $(\beta_{\rm QED},\ \beta_{\rm LogP})$ using the same procedure as above, and from Fig. 3, we see that we target multiple properties with varying fidelity by simply modulating the value of property-specific $\beta.$ Ultimately, increasing the $\beta$ for an individual property enables us to favor higher values of that property in multi-property alignment setting. In this case, we also do not regularize with the KL-divergence to the reference policy and again maintain sample diversity and validity (see Fig. S9 and Table S5)
3.1.2 Prompted molecular alignment on RDKit oracles
Inspired by the task of lead optimization in drug discovery efforts [21], we ask whether we can use ERA to train a molecular generator that can sample a molecule that is both similar to the prompt molecule and also exhibits some desired property. First, we fine-tune the pretrained molecular generator to enable prompted molecular generation (see Appendix E.3.2) and use this fine-tuned model as our reference policy for all prompted molecular alignment tasks. This reference policy disproportionately samples molecules that are identical (i.e. a Tanimoto similarity of 1.0) to the prompt molecule (see Fig. S10), so we carry out multi-property alignment on this reference policy to generate molecules that are similar—but not identical—to the prompt molecule and also have a high drug-likeness as measured by QED. Using ERA, we optimize the reference policy with a generated dataset $\mathcal{D}=\{(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)})\}_{i=1}^{N}$, where we sample four molecules for each prompt molecule from the reference policy and consider all possible preference pairs for a total of six preference pairs per prompt molecule (see Appendix E.2 for full details on the energy used).
We observe that the per-prompt average QED under the optimized policy for a given prompt is higher than the corresponding average under the reference policy (Fig. S10). Furthermore, we see that we are able to sample a diverse set of molecules that are chemically similar to the prompt molecule, and also chemically valid (see Fig. S11, Table S6). We repeat the experiment with a related objective of generating molecules similar to the prompt molecule with a high LogP instead and again observe that we increase the per-prompt average LogP under the optimized policy relative to the reference policy without degrading sample diversity and validity. For both of these experiments, we required regularization to the reference policy ($\gamma>0$). With no regularization, the aligned generator would almost exclusively sample sequences that were chemically invalid ($<25\%$ chemical validity). Finally, we note that the increases in QED and LogP in Fig. S10 are smaller relative to the increases in Fig. 2 because the samples are now conditioned to remain proximal to the prompt molecule, which restricts the chemical space that can be explored.
[…]
mean score IntDiv mean score IntDiv
ERA 0.996 $\pm$ 0.000 0.219 $\pm$ 0.002 0.987 $\pm$ 0.001 0.264 $\pm$ 0.005
GraphGA 0.919 $\pm$ 0.016 0.365 $\pm$ 0.024 0.875 $\pm$ 0.025 0.380 $\pm$ 0.015
JT-VAE 0.235 $\pm$ 0.083 0.770 $\pm$ 0.067 0.159 $\pm$ 0.040 0.781 $\pm$ 0.127
[…]
We additionally measure sample efficiency using the top-10 AUC metric [14], which is the area under the curve (AUC) of the mean property value of the top-10 performing molecules versus the number of oracle calls (see Fig. S13 and Table S7). We likewise only include novel, unique, and valid molecules in this analysis; any sampled molecule that is in ChemBL, that has already been sampled, or that is invalid is discarded and additionally does not count towards an oracle call as these are filtered out before oracle evaluation. We observe that we are able to generate novel and unique high-scoring molecules, with high sample efficiency especially in comparison to existing state-of-the-art methods. Ultimately, high sample efficiency is crucial in settings where evaluation is expensive, which will generally be true for most real-world chemical and biological tasks (e.g. wet-lab experiment). Finally, we also perform Glide Standard Precision [13] docking on the top-scoring molecules according to the oracles (score of 1.0) against their respective receptors. We observe that the diverse set of sampled molecules exhibit chemically plausable docked poses obtained from a physics-based docking approach (Fig. 4).
[…]
3.2 Directed evolution of proteins with ERA
We also consider the performance of ERA in a large-molecule setting, namely ML-guided directed evolution of proteins. Directed evolution campaigns aim to optimize a protein sequence toward some desired property of interest via iterative mutagenesis, library screening, and selection of best variants [41]. This has become a widely used methodology in protein engineering but comes with key limitations. The inherently iterative nature of directed evolution campaigns can lead to costly and time-consuming experimental campaigns, and meaningfully understanding the effects of protein mutations on protein activity can often be difficult. These challenges have led to the application of machine learning methods to more efficiently guide directed evolution campaigns [45] [36]. Given the success of ERA in guiding the optimization of small-molecules using a SMILES language representation, we examined whether ERA could be used to optimize large protein molecules using a protein language (i.e. primary sequence) representation.
There has been significant recent effort to design and train large protein language models (PLMs) [22, 17]. Furthermore, these models have demonstrated remarkable capabilities across a number of protein tasks [42, 36]. As such, we decided to use the state-of-the-art ESM3-1.4B [17] as our pretrained model, for which we carried out alignment using ERA. Despite the multimodal nature of ESM3, here, we only focus on generating primary-sequence-based representations of proteins.
We consider directed evolution of the $\beta$-subunit of tryptophansynthase (TrpB) from Thermotoga maritima, an enzyme that catalyzes tryptophan production [8]. Here, we seek to evolve the protein to increase its evolutionary fitness. In this work, we do not have access to experimental validation and so we evaluate the fitness of sequences using the computationally evaluable EVmutation score, an oracle that is predictive of a variant sequence’s performance relative to the parent sequence in its native function [18].
As in other directed evolution campaigns for the TrpB protein [44], we consider mutating four different sites to one of the 20 standard amino acids. We randomly sampled 512 mutated sequences, emulating a random mutagenesis experiment. Using ESM3-1.4B as our reference model, we carry out alignment using ERA with various $\beta=(0.1,1.0,10.0)$ and $\gamma=0.001$ and plot the results in Fig. 5 (see Appendix F for more details). We observe that with higher $\beta$, we are able to sample mutants with the highest possible EVmutation score in a single round of alignment. These results are promising for the application of ERA in directed evolution campaigns and future work will focus on the guidance of wet-lab directed evolution campaigns in conjunctions with multi-round, on-policy ERA.
4 Conclusions and Limitations
This paper introduces energy rank alignment, a simple and effective algorithm for policy optimization with an explicit reward model. We find that ERA is stable without extensive hyperparameter tuning, and sufficiently general to successfully align application-specific transformers for chemical search problems and protein language models. The algorithm exhibits strong performance with a variety of reward models, even ones with relatively weak signal. We analyze the minimizers of the ERA objective and find that they differ from the minimizers of popular policy alignment algorithms DPO and PPO in an important way: unlike PPO, the strength of regularization to the reference policy that we add is controlled by a parameter $\gamma$, while the entropy of the target distribution is independently tuned by a distinct parameter $\beta$. This means that we can avoid greedy policies by keeping $\beta$ small—amplifying fluctuations around the optimum of the reward model $(-U)$—while reducing the influence of the reference policy by taking $\gamma$ small. Our objective leads to easily interpretable sample-wise gradients which highlight the importance of a reward model relative to DPO in the sampled objective.
Limitations:
First, our approach requires a reward model, which can be difficult to train or design, especially for complex tasks. While we observed that ERA makes an appreciable impact even with weak supervision, this sort of proxy may not be available for more complex tasks. For example, optimizing small molecules for high binding affinity to a target protein would require expensive and noisy evaluations of a reward model, which likely limits the scope of molecular design to problems where the reward can be computed somewhat efficiently. A second limitation of our present work is that we do not train the molecular transformer to favor synthetic accessibility nor do we explicitly seek to obtain molecules that are easily synthesized experimentally. There are models that seek to evaluate synthesizability computationally that could be used in our rewards, which we plan to explore in future work [12].
Appendix A Code Availability
Appendix B Theoretical Analysis
To understand the ERA loss function and its connection to the entropy regularized objective (2), we first establish that the minimizers of (19) are of the form (5). We first define the notion of equivalence precisely.
Definition B.1
Lemma B.1
We prove Lemma B.1 in Appendix C and use this simple lemma to prove the following result.
Proposition B.2
from which we see that
By construction, $\pi(\boldsymbol{y}|\boldsymbol{x})$ does not depend on $\boldsymbol{y}^{\prime}$ so the prefactor must be purely a function of $\boldsymbol{x}$, which completes the proof, using Lemma B.1.
Appendix C Detailed Theoretical Analysis
Set-up, notation, and assumptions
Let $\mathcal{X}$ and $\mathcal{Y}$ be discrete spaces; each element of one of these spaces is a finite-length sequence of tokens within a fixed dictionary on which an autoregressive generative model is trained. The resulting models yield “policies”, which are conditional probability distributions $\pi(\cdot|\boldsymbol{x})\in\mathcal{P}(\mathcal{Y})$ for each $\boldsymbol{x}\in\mathcal{X}$. Throughout, we assume that our policies have full support on $\mathcal{Y}$ for each $\boldsymbol{x}$, meaning that $\inf_{\boldsymbol{y},\boldsymbol{x}\in\mathcal{Y}\times\mathcal{X}}\pi(\boldsymbol{y}|\boldsymbol{x})>0.$ Because the spaces are discrete, we make no strong restrictions on the regularity or coerciveness of the reward model $-U:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. The only requirement to ensure the existence of an optimal probability distribution is that $\sup_{\boldsymbol{x},\boldsymbol{y}\times\mathcal{X}\times\mathcal{Y}}|e^{-U(\boldsymbol{x},\boldsymbol{y})}|<+\infty$, which maintains full support of the distribution. Though it plays little role in theoretical analysis, we also denote by $\nu\in\mathcal{P}(\mathcal{X})$ the probability distribution over the prompts $\boldsymbol{x}.$
Goals of the analysis presented here
Definition C.1
Lemma C.1
If $g$ is such that
then the normalized policy $\pi_{g}^{\prime}$ is clearly defined by
Using Lemma  C.1 it is straightforward to prove the result in Proposition B.2. For completeness, we re-state that result here and refer the reader to Appendix B for the complete argument.
Proposition C.2
This proposition establishes that a policy minimizing the objective
$p_{\boldsymbol{\theta}}$ $:=\sigma\left(\log\frac{\pi_{\boldsymbol{\theta}}(\boldsymbol{y}|\boldsymbol{x})}{\pi_{\boldsymbol{\theta}}(\boldsymbol{y}^{\prime}|\boldsymbol{x})}\right)$
$p_{\gamma}$ $:=\sigma\left(\frac{\beta}{1+\gamma}\left[(U(\boldsymbol{x},\boldsymbol{y}^{\prime})-U(\boldsymbol{x},\boldsymbol{y}))+\beta^{-1}\gamma\log\frac{\pi_{\rm ref}(\boldsymbol{y}|\boldsymbol{x})}{\pi_{\rm ref}(\boldsymbol{y}^{\prime}|\boldsymbol{x})}\right]\right),$
has the form
C.1 Loss functions for $\pi_{\boldsymbol{\theta}}$:
Proximal Policy Optimization (PPO) optimizes an indirect, proximal objective to minimize an objective closely related to (2) (cf. Appendix C). Direct Preference Optimization (DPO) treats the negative reward function $U$ implicitly and directly maximizes the likelihood of $p(\boldsymbol{y}\succ\boldsymbol{y}^{\prime}|\boldsymbol{x})$. Our objectives differ from both approaches: like DPO, we directly optimize the policy using an explicit, gradient-based objective, but, in contrast, we use a reward function directly in our objective. The losses we build are thus amenable to both offline (samples from $\pi_{\rm ref}$) and online (samples from $\pi_{\boldsymbol{\theta}}$) policy alignment, as explained below. Choosing to optimize the objective online has been shown to have important consequences on performance [38], though we focus here on the setting where samples are drawn offline.
where
C.2 Gradients of $\mathcal{L}^{\textrm{ERA}}$.
@@ C.1 Loss functions for $\pi_{\boldsymbol{\theta}}$:
[…]
C.3 Comparison with PPO Objective
@@
Here, the KL-regularization corresponds to an energy shift, as in our objective, but there is no limit in which the ideal distribution $\pi\propto e^{-\beta U}$ is obtained for the PPO objective. This is in stark contrast to our approach, which recovers the ideal distribution as $\gamma\to 0.$ Furthermore, while our approach allows for a direct gradient-based optimization using (17), PPO is implemented using an actor-critic framework that is difficult to tune [29, 9]. Finally, we emphasize that for ERA in the $\gamma\to 0$, finite $\beta>0$, the distribution has positive entropy and is not manifestly mode-seeking; there can still be appreciable fluctuations in the output. Eliminating the effect of regularization in (28), on the other hand, requires taking $\beta/\gamma\to\infty$, which eliminates fluctuations in the distribution.
C.4 Comparison with DPO Objective
The DPO approach also seeks to optimize the objective (LABEL:eq:app_ppo_obj). The algorithm does so by first using (28) to define an implicit reward model by solving for the $U$ that reflects the observed preference probabilities. This elegant idea has had a significant impact and has already been deployed in state-of-the-art models [1]. In many cases, the observed preference probabilities will be sampled and only perhaps only one observation of $\boldsymbol{y}\succ\boldsymbol{y}^{\prime}$ will be available for each $\boldsymbol{x}$ in the dataset. When the preference dataset only has one observation $\boldsymbol{y}\succ\boldsymbol{y}^{\prime}$ per prompt $\boldsymbol{x}$, the optimal policy requires that
@@ C.4 Comparison with DPO Objective
which when $\pi_{\boldsymbol{\theta}}(\boldsymbol{y}^{\prime}|\boldsymbol{x})\to 0$, could lead to instability as $-\log\pi_{\boldsymbol{\theta}}(\boldsymbol{y}^{\prime}|\boldsymbol{x})\to\infty.$ On the other hand, the ERA gradients are scaled by the relative preference discrepancy,
@@
With ERA, we see that we are able to similarly sample high QED small-molecules with reasonably high chemical validity ($\sim$85%). While the validity does drop over subsequent checkpoints, it does not do so precipitously. Moreover, the ERA-based alignment had no regularization ($\gamma=0$), and in the main text, we document how having non-zero $\gamma$ can enable increases in chemical validity (Figure S10). Finally, we note that we did not extensively tune the hyperparameters for DPO, and it is possible that a different set of hyperparameters would elicit a more desired outcome; however, the lack of meaningful regularization in DPO [3] and its performance degradation in online metrics has been well-documented [10].
Appendix D ERA implementation
@@
Appendix E Details for molecular generator experiments
E.1 Pretraining details
In this work, we represent all molecules as SMILES strings and tokenize SMILES strings according to the approach in [34]. Our dataset consisted of all small-molecules from the ChEMBL database that were of length 500 tokens or less. Ultimately, this token limit filtered out approximately 0.1% of the small-molecules in the original ChEMBL dataset. The alphabet generated from this curated dataset consists of 324 tokens, which we augmented with start, stop, and padding tokens.
@@ E.1 Pretraining details
Our trained molecular generator consisted of just the encoder block of a standard multi-head attention transformer [39]. Finally, the model had 2 layers, 8 heads, and a width of 512. For pretraining, we used an Adam optimizer with a learning rate of $1.0*10^{-5}$. We emphasize that this pretrained generator samples molecules in an unprompted fashion; given just a start-of-sequence token, we can autoregressively generate a sequence of tokens. Moreover, it is possible that this sequence of tokens corresponds to a molecule that is not chemically valid, and we find that around 88% of all generated molecules are chemically valid. Lastly, we measure the diversity of the pretrained molecular generator by first generating 1500 molecules and then computing the Tanimoto similarity between every pair of molecules. We plot the distribution of all pairwise Tanimoto similarities from this sample and from all pariwise Tanimoto similarities from 1500 randomly sampled molecules from the original dataset in Fig. S7. We observe that we can generate molecules that are quite distinct (i.e. low Tanimoto similarity) in comparison with all other molecules.
@@
E.2 Chemical properties
We investigated aligning the molecule generator to several target chemical properties, which we detail below. All of the properties can be easily computed using either the RDKit package or the tdc [20] package. We list the energy function and parameters used for the corresponding energy functions for each of these properties in Table S2.
Tanimoto similarity is a measure of chemical and structural properties between two molecules and ranges from 0 to 1, where higher values correspond to more similar molecules [30]. Quantitative estimation of drug-likeness (QED) is evaluated by taking the geometric mean of a set of “desirability functions” for different molecular descriptors and also ranges continuously from values of 0 to 1 [6], where higher values correspond to more drug-like molecules. The octanol-water parition coefficient (Wildman-Crippen LogP) is a measure of hydrophobicity frequently employed in medicinal chemistry applications [43]. Molecules with more positive values are more hydrophobic (i.e. more soluble in octanol relative to water), whereas molecules with more negative values are more hydrophilic (i.e. more soluble in water relative to octanol). Molar refractivity is similarly calculated as a linear combination of atomic contributions, and is a positive number that serves as a measure for molecular size and polarizability [43]. A higher molar refractivity corresponds to larger and more polarizable molecules. Finally, ring count corresponds to the number of rings in a molecule.

### v3

@@ A.1 Loss functions for $\pi_{\boldsymbol{\theta}}$:
A.2 Gradients of $\mathcal{L}^{\textrm{ERA}}$.
@@
This calculation highlights one key difference between the approach we use and DPO. When the data only contains one observation of $\boldsymbol{y}\succ\boldsymbol{y}^{\prime}$ for a given $\boldsymbol{x},$ the DPO objective’s implicit reward model assigns zero probability to $\boldsymbol{y}^{\prime}\succ\boldsymbol{y}$. This pushes the policy towards extremal values, which can lead to undesired behavior, as discussed in Azar et al. [2023]. In our formulation, this behavior occurs only when the reward model assigns an energy of $\pm\infty$, which is prohibited by construction in most tasks. We further discuss differences between ERA and DPO in Appendix A.4.
A.3 Comparison with PPO Objective
@@
Here, the KL-regularization corresponds to an energy shift, as in our objective, but there is no limit in which the ideal distribution $\pi\propto e^{-\beta U}$ is obtained for the PPO objective. This is in stark contrast to our approach, which recovers the ideal distribution as $\gamma\to 0.$ Furthermore, while our approach allows for a direct gradient-based optimization using (17), PPO is implemented using an actor-critic framework that is difficult to tune Rafailov et al. [2023], Casper et al. [2023]. Finally, we emphasize that for ERA in the $\gamma\to 0$, finite $\beta>0$, the distribution has positive entropy and is not manifestly mode-seeking; there can still be appreciable fluctuations in the output. Eliminating the effect of regularization in (28), on the other hand, requires taking $\beta/\gamma\to\infty$, which eliminates fluctuations in the distribution.
A.4 Comparison with DPO Objective
The DPO approach also seeks to optimize the objective (LABEL:eq:app_ppo_obj). The algorithm does so by first using (28) to define an implicit reward model by solving for the $U$ that reflects the observed preference probabilities. This elegant idea has had a significant impact and has already been deployed in state-of-the-art models AI@Meta [2024]. In many cases, the observed preference probabilities will be sampled and only perhaps only one observation of $\boldsymbol{y}\succ\boldsymbol{y}^{\prime}$ will be available for each $\boldsymbol{x}$ in the dataset. When the preference dataset only has one observation $\boldsymbol{y}\succ\boldsymbol{y}^{\prime}$ per prompt $\boldsymbol{x}$, the optimal policy requires that
@@ A.4 Comparison with DPO Objective
which when $\pi_{\boldsymbol{\theta}}(\boldsymbol{y}^{\prime}|\boldsymbol{x})\to 0$, could lead to instability as $\log\pi_{\boldsymbol{\theta}}(\boldsymbol{y}^{\prime}|\boldsymbol{x})\to\infty.$ On the other hand, the ERA gradients are scaled by the relative preference discrepancy,
@@
With ERA, we see that we are able to similarly sample high QED small-molecules with reasonably high chemical validity ($\sim$85%). While the validity does drop over subsequent checkpoints, it does not do so precipitously. Moreover, the ERA-based alignment had no regularization ($\gamma=0$), and in the main text, we document how having non-zero $\gamma$ can enable increases in chemical validity (Figure S10). Finally, we note that we did not extensively tune the hyperparameters for DPO, and it is possible that a different set of hyperparameters would elicit a more desired outcome; however, the lack of meaningful regularization in DPO Azar et al. [2023] and its performance degradation in online metrics has been well-documented Chen et al. [2024].
Appendix B ERA implementation
@@
Appendix C Details for molecular generator experiments
C.1 Pretraining details
In this work, we represent all molecules as SMILES strings and tokenize SMILES strings according to the approach in Schwaller et al. [2018]. Our dataset consisted of all small-molecules from the ChEMBL database that were of length 500 tokens or less. Ultimately, this token limit filtered out approximately 0.1% of the small-molecules in the original ChEMBL dataset. The alphabet generated from this curated dataset consists of 324 tokens, which we augmented with start, stop, and padding tokens.
@@ C.1 Pretraining details
Our trained molecular generator consisted of just the encoder block of a standard multi-head attention transformer Vaswani et al. [2017]. Finally, the model had 2 layers, 8 heads, and a width of 512. For pretraining, we used an Adam optimizer with a learning rate of $1.0*10^{-5}$. We emphasize that this pretrained generator samples molecules in an unprompted fashion; given just a start-of-sequence token, we can autoregressively generate a sequence of tokens. Moreover, it is possible that this sequence of tokens corresponds to a molecule that is not chemically valid, and we find that around 88% of all generated molecules are chemically valid. Lastly, we measure the diversity of the pretrained molecular generator by first generating 1500 molecules and then computing the Tanimoto similarity between every pair of molecules. We plot the distribution of all pairwise Tanimoto similarities from this sample and from all pariwise Tanimoto similarities from 1500 randomly sampled molecules from the original dataset in Fig. S7. We observe that we can generate molecules that are quite distinct (i.e. low Tanimoto similarity) in comparison with all other molecules.
@@
C.2 Chemical properties
We investigated aligning the molecule generator to several target chemical properties, which we detail below. All of the properties can be easily computed using either the RDKit package or the tdc Huang et al. [2021] package. We list the energy function and parameters used for the corresponding energy functions for each of these properties in Table S2.
Tanimoto similarity is a measure of chemical and structural properties between two molecules and ranges from 0 to 1, where higher values correspond to more similar molecules Rogers and Tanimoto [1960]. Quantitative estimation of drug-likeness (QED) is evaluated by taking the geometric mean of a set of “desirability functions” for different molecular descriptors and also ranges continuously from values of 0 to 1 Bickerton et al. [2012], where higher values correspond to more drug-like molecules. The octanol-water parition coefficient (Wildman-Crippen LogP) is a measure of hydrophobicity frequently employed in medicinal chemistry applications Wildman and Crippen [1999]. Molecules with more positive values are more hydrophobic (i.e. more soluble in octanol relative to water), whereas molecules with more negative values are more hydrophilic (i.e. more soluble in water relative to octanol). Molar refractivity is similarly calculated as a linear combination of atomic contributions, and is a positive number that serves as a measure for molecular size and polarizability Wildman and Crippen [1999]. A higher molar refractivity corresponds to larger and more polarizable molecules. Finally, ring count corresponds to the number of rings in a molecule.